OpenStudy (anonymous):

A number is chosen at random from the first 100 positive integers. Find the probability that the number is either even or a perfect square. 11/20 1/20

OpenStudy (anonymous):

@trindawg

OpenStudy (anonymous):

\[P \left( A \cup B \right)=P(A)+P(B)-P(A~intersection ~B)\] A= event even no. B=event perfect square. A={2,4,6,8,....,98,100},n(A)=50 B={1,4,9,16,25,36,49,64,81,100},n(B)=10 A intersection B={4,16,36,64,100},n(A intersection B )=5 \[P(A \cup B )=?\]

OpenStudy (anonymous):

\[P(A \cup B)=\frac{ 50 }{ 100 }+\frac{ 10 }{ 100 }-\frac{ 5 }{ 100 }=?\]

OpenStudy (xapproachesinfinity):

1/2+1/10-1/20=?

OpenStudy (xapproachesinfinity):

the ways @surjithayer wrote is better :)

OpenStudy (anonymous):

11/20?

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

thnx

OpenStudy (xapproachesinfinity):

seems to me you just guessed heheh